OpenStudy (anonymous):

Given the exponential equation 3x = 243, what is the logarithmic form of the equation in base 10?

OpenStudy (idealist10):

3^5=243

OpenStudy (ikram002p):

u mean \(3^x=243\)

OpenStudy (ikram002p):

just solve for x :P

OpenStudy (idealist10):

x=ln(243)/ln(3)=5

OpenStudy (anonymous):

ooops im sorry i forgot to say its just setting it up... x = log base 10 of 243, all over log base 10 of 3 x = log base 10 of 3, all over log base 10 of 243 x = log base 2 of 3, all over log base 2 of 243 x = log base 2 of 243 all over log base 2 of 3

OpenStudy (anonymous):

@ganeshie8

OpenStudy (thebluepilot):

Did you ever get the answer @Avenged7x ?